OpenStudy (anonymous):

solve the following system of equations: y=x+4 2x+y=13 (1 point) A-(3,7) B-(5,8) C-(9,5) D-(7,3) (PLEASE HELP)

jimthompson5910 (jim_thompson5910):

2x+y=13 2x+x+4=13 ... replace y with x+4, since y = x+4 now solve for x